Exercise 6-6

Presented below are three unrelated situations.

Stellar Company recently signed a lease for a new office building, for a lease period of 11 years. Under the lease agreement, a security deposit of $14,200 is made, with the deposit to be returned at the expiration of the lease, with interest compounded at 5% per year.

Click here to view factor tables

What amount will the company receive at the time the lease expires? (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)
The company will receive $enter a dollar amount rounded to 0 decimal places

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

LINK TO VIDEO

Pearl Corporation, having recently issued a $20,066,900, 15-year bond issue, is committed to make annual sinking fund deposits of $623,400. The deposits are made on the last day of each year and yield a return of 10%.

Click here to view factor tables

Will the fund at the end of 15 years be sufficient to retire the bonds?
Future value of an ordinary annuity $enter a dollar amount rounded to 0 decimal places

Will funds be sufficient?

select between Yes and No

NoYes

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

LINK TO VIDEO

Pearl Corporation, having recently issued a $20,066,900, 15-year bond issue, is committed to make annual sinking fund deposits of $623,400. The deposits are made on the last day of each year and yield a return of 10%.

Click here to view factor tables

Determine the amount of deficiency. (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)
Deficiency $enter a dollar amount rounded to 0 decimal places

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

LINK TO VIDEO

Under the terms of his salary agreement, president Newman Walters has an option of receiving either an immediate bonus of $55,000, or a deferred bonus of $70,000 payable in 10 years.

Click here to view factor tables

Ignoring tax considerations and assuming a relevant interest rate of 4%, which form of settlement should Walters accept?
Present value of deferred bonus $enter a dollar amount rounded to 0 decimal places

select an option

Accept Deffered Bonus in 10 yearsAccept Bonus Now

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

LINK TO VIDEO
Answer #1

a)Stellar Company has paid a security deposit of $14,200 , with the deposit to be returned at the expiration of the lease of 11 years, with interest compounded at 5% per year.

Security deposit that will be received by Stellar company will be 14200*(1.05)^11=24,286.82

b)Pearl Corporation, having recently issued a $20,066,900, 15-year bond issue, is committed to make annual sinking fund deposits of $623,400. The deposits are made on the last day of each year and yield a return of 10%.

Value of sinking fund at the end of 15 years= 623400*future value annuity factor

=623400*31.772

=19,806,965

The amount at the end of fund wil not be sufficient

c) amount of deficiency will be 15 year bonds amount- sinking fund value at the time of redemption

=20,066,900-19,806,965

=259,935

d)Under the terms of his salary agreement, president Newman Walters has an option of receiving either an immediate bonus of $55,000, or a deferred bonus of $70,000 payable in 10 years.

Present value of deferred bonus discounted at 4%=70000/(1.04)^10

=47289.5

Walters should agree for immediate payment of bonus as this bonus has higher value when compared to deferred bonus.